# A Modified Center-of-Mass Conservation Law in Finite-Domain Simulations of the Zakharov–Kuznetsov Equation

### v1

@@ I Introduction
In this context, it is natural to ask if numerical simulations in a finite-domain also hold the conservation laws defined in the whole space. To address this question, we focus on the two-dimensional Zakharov–Kuznetsov (ZK) equation, which serves as a natural higher-dimensional generalization of the Korteweg–de Vries (KdV) equation. Although the ZK equation admits several conservation laws on the whole space, we show that one of them—the center-of-mass-type invariant—exhibits a systematic variation in standard finite-domain numerical simulations. This variation is not a numerical artifact but rather a consequence of boundary-flux contributions that are absent in the infinite-domain theory. By explicitly accounting for these contributions, we derive a modified conserved quantity that restores consistency between analytical theory and numerical computation.
A classical and influential example of a nonlinear dispersive PDE is the KdV equation Zabusky and Kruskal (1965). It is a prototypical integrable system, possessing an infinite hierarchy of conservation laws that support stable soliton solutions Miura et al. (1968). The structure of conserved quantities is fundamental both for understanding qualitative features of solutions and for constructing accurate and robust numerical schemes.
In contrast to fully integrable systems like the KdV equation, many multi-dimensional dispersive PDEs permit only a finite set of conservation laws and are often referred to as quasi-integrable systems Ferreira et al. (2013). The ZK equation belongs to this class. It arises as an asymptotic model for ion–acoustic waves in magnetized plasma Sagdeev and Galeev (1969); Zakharov and Kuznetsov (1974) and has been extensively studied from both analytical and numerical perspectives Kuznetsov et al. (1986); Iwasaki et al. (1990); Faminskii (1995); de Bouard (1996); Munro and Parkes (1999); Linares and Pastor (2009); Johnson (2010); Linares and Pastor (2011); Farah et al. (2012); Ribaud and Vento (2012); Grø”unrock and Herr (2013); Pilod (2014); Ponce (2015); Côté et al. (2016); Klein et al. (2021); Faminskii (2022). The ZK equation admits three standard scalar conserved quantities—mass, momentum, and energy—denoted by $I_{i}$ $(i=1,2,3)$, as well as a vector-valued quantity ${\mbox{$I$}}_{4}$ associated with the center of mass Zakharov and Kuznetsov (1974). The scalar invariants have been discussed extensively in the literature, and their numerical preservation has been examined in a variety of computational studies Bridges and Reich (2001); Nishiyama et al. (2012); Xu et al. (2012); Klein et al. (2023). In contrast, comparatively little attention has been paid to the vector-valued quantity that characterizes the motion of the center of mass.
@@
Since explicit solutions are available only in special cases, numerical simulation plays a central role in the study of the ZK equation. Fourier (pseudo-)spectral discretizations Gottlieb and Orszag (1977); Trefethen (2000), often combined with exponential time-differencing Runge–Kutta schemes Cox and Matthews (2002); Kassam and Trefethen (2005), have become standard tools due to their excellent accuracy and stability for dispersive equations Chan and Kerkhoven (1985); Gavete et al. (2008); Chen et al. (2011); Zavalani (2014); Klein and Peter (2014); Klein et al. (2021). These methods are known to reproduce the standard scalar conserved quantities with high accuracy Bridges and Reich (2001); Chen et al. (2011). However, the numerical behavior of the center-of-mass quantity, particularly its sensitivity to finite-domain truncation and boundary effects, has not been systematically investigated.
In this paper, we perform high-precision numerical simulations of the two-dimensional ZK equation on a finite domain. The standard scalar invariants, mass, momentum, and energy, are preserved to high accuracy. The $y$-component of the center-of-mass quantity is likewise well conserved. In contrast, the $x$-component exhibits a clear and systematic temporal drift, despite being an exact conserved quantity of the equation. We show that this deviation originates from boundary-flux contributions that arise in the finite-domain formulation. Motivated by this observation, we introduce a modified center-of-mass conserved quantity that explicitly subtracts the analytically derived boundary contribution and demonstrate its conservation in numerical experiments.
@@ II Zakharov–Kuznetsov Equation
$\frac{\partial u}{\partial t}+\alpha\,u\frac{\partial u}{\partial x}+\frac{\partial}{\partial x}(\Delta u)=0,$ (1)
where the Laplacian is defined as $\Delta=\partial_{x}^{2}+\partial_{y}^{2}$, and the parameter $\alpha$ controls the strength of the nonlinear term, commonly taken to be $\alpha=2$. This parameter can be absorbed into the amplitude of the solution. In fact, by introducing the rescaling $u=\beta v$, Eq. (1) transforms into
$v_{t}+(\alpha\beta)\,vv_{x}+\partial_{x}\Delta v=0.$
@@
$-c\,U_{\tilde{x}}+\alpha\,U\,U_{\tilde{x}}+\partial_{\tilde{x}}\Delta U=0.$
@@
$\Delta U=c\,U-\frac{\alpha}{2}U^{2}.$ (2)
A one-dimensional traveling solitary-wave solution within our convention $\alpha=12$ is given by
@@ III.1 Numerical Simulations
We set a doubly periodic computational domain $(x,y)\in[-L_{x},L_{x}]\times[-L_{y},L_{y}]$ with $(L_{x},L_{y})=(32.0,16.0)$. The numbers of grid points are chosen as $(N_{x},N_{y})=(1024,512)$. Time integration is performed over the interval $t\in[0,10]$ using a time step $\Delta t=2.5\times 10^{-4}$. Spatial discretization is carried out using a Fourier pseudo-spectral method, combined with an exponential time-differencing fourth-order Runge–Kutta (ETDRK4) scheme for time integration. These parameters are chosen to minimize numerical dissipation while ensuring adequate spatial resolution throughout the simulation.
@@ III.2 Analytical Study of the Boundary-Induced Variation in $I_{4x}$
$u_{t}+\partial_{x}Q=0,\qquad Q:=6u^{2}+\Delta u.$ (8)
@@ Acknowledgments
The authors thank Luiz Agostinho Ferreira, Atsushi Nakamula, Ryu Sasaki, Kiori Obuse and Kouichi Toda for their fruitful discussions and valuable comments. N. S. is supported in part by JSPS KAKENHI Grant Number JP23K02794.
@@ Appendix A Boundary Effects on $I_{4x}$
$\iint_{D}Q\,dxdy$ $=\iint_{D}\Delta u\,dxdy+6\iint_{D}u^{2}\,dxdy,$
$\iint_{D}\Delta u\,dxdy$ $=\int_{-L_{y}}^{L_{y}}[\,u_{x}(t,L_{x},y)-u_{x}(t,-L_{x},y)\,]dy$

### v2

@@ I Introduction
In this context, it is natural to ask whether finite-domain simulations also satisfy conservation laws that are derived on the whole space. To address this question, we consider the two-dimensional Zakharov–Kuznetsov (ZK) equation, a natural higher-dimensional generalization of the Korteweg–de Vries (KdV) equation. While the ZK equation admits several conserved quantities on the whole space, their behavior under finite-domain truncation, especially for center-of-mass-type invariants, is far less understood.
For context, a classical and influential example of a nonlinear dispersive PDE is the KdV equation Zabusky and Kruskal (1965). It is a prototypical integrable system, possessing an infinite hierarchy of conservation laws that support stable soliton solutions Miura et al. (1968). The structure of conserved quantities is fundamental both for understanding qualitative features of solutions and for constructing accurate and robust numerical schemes. Moreover, conserved quantities underpin variational and stability theories for Hamiltonian dispersive systems (see, e.g., Zakharov and Kuznetsov (2012)), and in the ZK case the linear stability of ground-state solitons has been rigorously linked to a Vakhitov–Kolokolov-type condition formulated via the conserved momentum Kuznetsov (2018).
In contrast to fully integrable systems like the KdV equation, many multi-dimensional dispersive PDEs permit only a finite set of conservation laws and are often referred to as quasi-integrable systems Ferreira et al. (2013). The ZK equation belongs to this class. It arises as an asymptotic model for ion–acoustic waves in magnetized plasma Sagdeev and Galeev (1969); Zakharov and Kuznetsov (1974) and has been extensively studied from both analytical and numerical perspectives Kuznetsov et al. (1986); Iwasaki et al. (1990); Faminskii (1995); de Bouard (1996); Munro and Parkes (1999); Linares and Pastor (2009); Johnson (2010); Linares and Pastor (2011); Farah et al. (2012); Ribaud and Vento (2012); Grø”unrock and Herr (2013); Pilod (2014-2015); Ponce (2015); Côté et al. (2016); Klein et al. (2021); Faminskii (2022). The ZK equation admits three standard scalar conserved quantities—mass, momentum, and energy—denoted by $I_{i}$ $(i=1,2,3)$, as well as a vector-valued quantity ${\mbox{$I$}}_{4}$ associated with the center of mass Zakharov and Kuznetsov (1974). The scalar invariants have been discussed extensively in the literature, and their numerical preservation has been examined in a variety of computational studies Bridges and Reich (2001); Nishiyama et al. (2012); Xu et al. (2012); Klein et al. (2023). In contrast, comparatively little attention has been paid to the vector-valued quantity that characterizes the motion of the center of mass.
@@
Since the explicit solutions are available only in special cases, numerical simulation plays a central role in the study of the ZK equation. Fourier (pseudo-)spectral discretizations Gottlieb and Orszag (1977); Trefethen (2000), often combined with exponential time-differencing Runge–Kutta schemes Cox and Matthews (2002); Kassam and Trefethen (2005), have become standard tools due to their excellent accuracy and stability for dispersive equations Chan and Kerkhoven (1985); Gavete et al. (2008); Chen et al. (2011); Zavalani (2014); Klein and Peter (2014); Klein et al. (2021). These methods are known to reproduce the standard scalar conserved quantities with high accuracy Bridges and Reich (2001); Chen et al. (2011). However, the numerical behavior of the center-of-mass quantity, particularly its sensitivity to finite-domain truncation and boundary effects, has not been systematically investigated.
In this paper, we introduce a modified center-of-mass conserved quantity that remains well conserved in finite-domain simulations. We perform high-precision numerical simulations of the two-dimensional ZK equation on a finite domain. The standard scalar conserved quantities—mass, momentum, and energy—are preserved to high accuracy. The $y$-component of the center-of-mass quantity is likewise well conserved. In contrast, the $x$-component exhibits a clear and systematic temporal drift, despite being an exact conserved quantity of the equation. We show that this deviation originates from boundary-flux contributions that arise in the finite-domain formulation. The modified quantity is therefore defined by explicitly subtracting the analytically derived boundary contribution, and we demonstrate its conservation in numerical experiments.
@@ II Zakharov–Kuznetsov Equation
$\frac{\partial u}{\partial t}+\alpha\,u\frac{\partial u}{\partial x}+\frac{\partial}{\partial x}(\nabla^{2}u)=0,$ (1)
where the Laplacian is defined as $\nabla^{2}=\partial_{x}^{2}+\partial_{y}^{2}$, and the parameter $\alpha$ controls the strength of the nonlinear term, commonly taken to be $\alpha=2$. This parameter can be absorbed into the amplitude of the solution. In fact, by introducing the rescaling $u=\beta v$, Eq. (1) transforms into
$v_{t}+(\alpha\beta)\,vv_{x}+\partial_{x}\nabla^{2}v=0.$
@@
$-c\,U_{\tilde{x}}+\alpha\,U\,U_{\tilde{x}}+\partial_{\tilde{x}}\nabla^{2}U=0.$
@@
$\nabla^{2}U=c\,U-\frac{\alpha}{2}U^{2}.$ (2)
A steady progressive solitary-wave solution within our convention $\alpha=12$ is given by
@@ III.1 Numerical Simulations
We set a doubly periodic computational domain $(x,y)\in[-L_{x},L_{x}]\times[-L_{y},L_{y}]$ with $(L_{x},L_{y})=(32,16)$. The numbers of grid points are chosen as $(N_{x},N_{y})=(1024,512)$. Time integration is performed over the interval $t\in[0,10]$ using a time step $\Delta t=2.5\times 10^{-4}$. Spatial discretization is carried out using a Fourier pseudo-spectral method, combined with an exponential time-differencing fourth-order Runge–Kutta (ETDRK4) scheme for time integration. These parameters are chosen to minimize numerical dissipation while ensuring adequate spatial resolution throughout the simulation.
@@ III.2 Analytical Study of the Boundary-Induced Variation in $I_{4x}$
$u_{t}+\partial_{x}Q=0,\qquad Q:=6u^{2}+\nabla^{2}u.$ (8)
@@ V Summary
Moreover, conservation-law-based diagnostics have also played an important role in stability considerations (e.g., Kuznetsov (2018)). In this context, the (modified) conserved quantity for center of mass, ${\mbox{$I$}}_{4}$, provides a complementary and practically useful diagnostic: it offers a robust way to separate genuine shape deformation from neutral translational dynamics in finite-domain simulations. This feature is particularly valuable in multi-pulse interactions, where accurate centroid tracking facilitates a systematic assessment of collision-induced displacement and phase shifts, as well as weak radiative effects, even when individual pulses partially overlap.
@@ Acknowledgments
The authors thank Luiz Agostinho Ferreira, Atsushi Nakamula, Ryu Sasaki, Kiori Obuse and Kouichi Toda for fruitful discussions and valuable comments. The authors are grateful to E. A. Kuznetsov for helpful comments and suggestions. N. S. is supported in part by JSPS KAKENHI Grant Number JP23K02794.
@@ Appendix A Boundary Effects on $I_{4x}$
$\iint_{D}Q\,dxdy$ $=\iint_{D}\nabla^{2}u\,dxdy+6\iint_{D}u^{2}\,dxdy,$
$\iint_{D}\nabla^{2}u\,dxdy$ $=\int_{-L_{y}}^{L_{y}}[\,u_{x}(t,L_{x},y)-u_{x}(t,-L_{x},y)\,]dy$